# A physics-informed neural network for modeling fracture without gradient damage: formulation, application, and assessment

### v1

@@ 4.1 Asymmetric double-edge notched plate
A specimen geometry often used to test the predictive capabilities of damage and fracture models in terms of crack path is a plate with asymmetric double-edge notches subjected to tension under plane-strain conditions (Ambati et al., 2016, Simoes and Martínez-Pañeda, 2021, Han et al., 2022). The loading schematic and specimen geometry considered are shown in Figure 2(A). The crack path predictions obtained using FEM with and without gradient damage for a mesh not aligned with the anticipated crack path are shown in Figure 2(B). As expected, the results from FEM without gradient damage solution are nonphysical. The significant mesh-dependency of FEM without gradient damage is shown through comparison of the predicted crack paths for two different meshes in Figure 2(C). The nonphysical crack paths are dissimilar for the two meshes. In FEM without gradient damage, upon failure of an element, the subsequent damage growth depends on the local mesh structure around that element, resulting in the solution’s strong mesh dependency.
For the collocation grid in Figure 3, the PINN was applied incrementally [referred to as PINN (incremental) hereon]. Specifically, the time step size was chosen to be 0.0025 s with the first increment at s. Training for each time increment was performed for 50000 epochs. The time interval from the initiation of damage to the complete rupture ( = 0.09 s) is small compared to the time to the initiation of damage (FEM with gradient damage: 0.477 s, PINN (incremental): 0.485 s), highlighting the rapid damage growth upon initiation built into the PINN formulation. The crack path predictions from the PINN (incremental) and the baseline FEM with gradient damage solution show good agreement, as shown in Figure 3.
@@ 4.2 Single-edge notched plate with a hole
A single-edge notched plate with a hole under plane-strain conditions is a specimen geometry utilized to validate damage and fracture models (Ambati et al., 2015, Sargado et al., 2021, Muixí et al., 2020). The loading schematic and specimen geometry considered are shown in Figure 4(A). The crack path predictions obtained using FEM with and without gradient damage for a mesh not aligned with the anticipated crack path are shown in Figure 4(B). The FEM without gradient damage solution is nonphysical. Figure 4(C) shows the physical solution using FEM without gradient damage for a mesh aligned with the anticipated crack path, highlighting the method’s mesh dependency.
PINN (incremental) is applied to the collocation grid in Figure 5(B) with time step size = 0.01 s, the first time increment at s, and 100000 epochs training for each time increment. The progression of damage at four different stages and the crack path in the PINN (incremental) solution agree with the FEM with gradient damage solution as seen in Figure 5.
We also consider a variation in this geometry where the hole is moved further away from the notch. Specifically, only the separation of the hole center from the notch centerline in Figure 4(A) is increased by 0.2. The crack path in the FEM with gradient damage solution changes significantly as seen in Figure 6. The PINN (incremental) [first increment at s, time step size = 0.01 s, 100000 epochs training for each time increment] captures the crack path variation with the defect positioning accurately as shown in Figure 6.
@@ 4.3 Plate with four randomly distributed holes
We now consider a relatively complex geometry consisting of four holes with different diameters randomly distributed on the 1x1 plate domain. The loading schematic, specimen geometry and finite element mesh considered are shown in Figure 7(A). The crack path prediction obtained using FEM with gradient damage is shown in Figure 7(B).
The PINN (incremental) is applied to the collocation grid in Figure 8(B) with time step size = 0.0025 s, the first time increment at s, and 200000 epochs training for each time increment. Good agreement of the damage progression at four different stages and the crack path from the PINN (incremental) solution with the baseline FEM with gradient damage solution can be seen in Figure 8.
@@ 5 Assessment of the PINN performance for fracture
Previously, we demonstrated the success of PINN without gradient damage formulation in predicting the fracture path for a variety of problems. The performance of any new and emerging computational method has to be evaluated to guide the choice of underlying parameters for future applications. Towards this, we investigate the effect of a few key neural network parameters on the PINN’s prediction performance. The PINN solutions in this subsection for the asymmetric double-edge notched plate geometry are obtained using a relatively computationally inexpensive single time increment approach [referred to as PINN (one-time step) hereon]. The crack paths are obtained through PINN (one-time step) solutions at s with 200000 epochs of training. relative errors and losses are evaluated through PINN (one-time step) application at s (before damage initiation) with 200000 epochs training. The state-dependent nature of damage progression implies PINN (incremental) fracture solutions will have higher accuracy than the computationally cheaper PINN (one-time step) solutions. However, the PINN (one-time step) solution provides crack paths with reasonable accuracy. This is demonstrated through a comparison of crack path predictions from the two methods for an asymmetric double-edge notched plate geometry, as shown in Appendix Figure A1. Hence, PINN (one-time step) can be used to study and obtain computationally optimum neural network parameters and architecture. First, variations in the collocation point distribution (approximately constant grid density) are considered. Further, the number of layers, the number of neurons in each layer, and the grid density (in a fixed grid structure) are systematically varied. The base number of layers, number of neurons in each layer, and grid density are the values used in Section 4, i.e., , , and , respectively. We also compare the computation time for the PINN using the base network and base grid density with those for FEM without and with gradient damage. The effects of simple pre-training and output normalization strategies on the solutions are also explored.
@@
relative error ( - displacement, - stress, UFE - undamaged free energy) and loss (total, - balance of linear momentum, - boundary conditions, - constitutive equations) evolution with number of epochs during training [PINN (one-time step), =0.45 s, 200000 epochs, grid from Figure 3] for the asymmetric double-edge notched plate geometry is shown in Figure 9. The errors and losses decrease monotonically with the number of epochs in this range. Corresponding results for the single-edge notched plate with a hole and the plate with four randomly distributed holes geometries are presented in the Appendix.
@@ 5.3 Effect of the number of layers, number of neurons in each layer, and grid density
The three different numbers of layers, number of neurons in each layer, and grid density considered were 3, 5, 8; 10, 30, 50; and , , , respectively. A numerical Factorial Design of Experiments (DOE) was performed to understand the main, simple, and interacting effects of the three hyperparameters. The complete results for the relative errors and the crack paths for the DOE study are presented in the Appendix. Figures 4(A), (B), (C) and (D) show the % errors for a few cases, highlighting the effect of (i) number of layers ( grid density), (ii) number of neurons in each layer ( grid density),(iii) grid density (5 layers), and (iv) grid density (30 neurons in each layer), respectively. All three factors have main effects. Significant interaction effects are observed between the number of layers and the number of neurons, as well as between the grid density and the number of layers. The PINN performance is optimized when the base values in Section 4 are used, i.e., 5 layers, 30 neurons in each layer, and grid density. Underfitting or overfitting occurs when the neural network parameters are not optimal.
5.4 Computation time: PINN, FEM without and with gradient damage
The computation time using the base network (5 layers each with 30 neurons) and the base grid density () for 1 epoch of PINN training remained approximately constant for all numerical problems considered and was equal to 0.06 s (GPU accelerated, QuadroRTX, CCV@Brown), 0.18 s (GPU accelerated, NVIDIA T600 & Intel Core i7-10700 Processor, Dell Precision 3650 Tower), and 0.24 s (not GPU accelerated, CCV@Brown). FEM without and with gradient damage (both implemented in ABAQUS/Explicit (2018)) were parallelized using 32 processors on a system with an Intel Xeon Gold 6326 CPU. For the finite element meshes considered in this work ( 10000 elements), the approximate computation time for an increment with a time step size of approximately s was 0.003 s for FEM with gradient damage.
PINN (incremental) in general requires more computation time than FEM. The complexity of implementing gradient damage models in FEM and the associated implementation time costs is an important consideration. Some reduction in computation time for PINN (incremental) can be achieved when the material response is rate-independent. As done in this work for an elastomer with hyperelastic response, the solution evaluation in PINN (incremental) for a rate-independent material can be started at any time instant before damage initiation. The computation time for PINN (one-time step), however, can be comparable to FEM.
@@ 5.5 Effect of pre-training
Pre-training has been shown in the literature to improve the accuracy and convergence of PINNs (Guo et al., 2023, Wang et al., 2025). We investigate the effect of one pre-training strategy. Specifically, the domain is seeded with collocation points as shown in Figure 12(A) with the same material properties and boundary conditions discussed in Section 4. The trained network parameters from stretching this unit square (=0.45 s in one time step, 200000 epochs) are used to initialize the PINN for application to the asymmetric double-edge notch plate geometry. Comparison between the PINN (one-time step) results (at =0.45 s [before damage initiation] and 0.65 s [after damage initiation], 200000 epochs) with and without pretraining in Figures 12(B) and (C) shows the simple pre-training associated convergence and accuracy enhancement.
@@ 5.6 Effect of PINN output normalization
Normalization of PINN outputs is another technique that has been shown in the literature to enhance accuracy and convergence (Xu et al., 2024, Rasht-Behesht et al., 2022). We consider a simple normalization scheme in the form of fixed normalization numbers that ensures physical consistency. Specifically, are normalized by , respectively before calculating the loss functions in equations (3.11), (3.13), (3.14). Figure 13 shows the comparison between the PINN (one-time step) results (at =0.45 s [before damage initiation] - , , , 200000 epochs and 0.65 s [after damage initiation] - , , ) for the asymmetric double-edge notched plate geometry with and without output normalization. Faster convergence and increased accuracy are achieved with this simple normalization scheme.
@@ 6 Conclusions
The important and evolving field of fracture modeling commonly uses the finite element method (FEM). FEM without gradient damage is relatively easier to implement. However, the corresponding results are significantly mesh-dependent and often inaccurate. FEM with gradient damage enables mesh-independent predictions at the cost of increased mathematical and numerical implementation complexities. Specifically, additional spatial gradient calculations and complex user-defined element and user subroutine development are necessary. Physics-informed neural networks (PINNs) are recently emerging as a computational method that can be potentially used to solve solid mechanics problems. The limited studies on PINNs for damage and fracture modeling in the literature are confined to the small-strain regime and use the gradient damage formulation by default. Due to the meshless nature of PINNs, we posed the question whether we can correctly solve fracture problems without needing the gradient damage formulation. We propose a PINN without gradient damage to model large deformation fracture in elastomers. The PINN does not require training data and utilizes the collocation method. The capability of the PINN to accurately predict the crack path has been validated by studying a variety of benchmark defect configuration problems, where we used accurate solutions from FEM with gradient damage and highly refined mesh. We also conducted an assessment of the PINN as a fracture modeling method by systematically varying a few key parameters to guide future design of PINNs for fracture problems. The PINN crack path predictions are approximately insensitive to the collocation point distribution. The proposed modeling framework and neural network architecture could motivate the application of PINNs without gradient damage for modeling fracture in a broad class of materials.
The relatively new nature of PINNs as a numerical method implies that the solution quality depends on a range of factors, and problem-dependent insights are necessary for the successful application of PINNs. However, the relative ease of mathematical formulation, numerical implementation, and mesh-free nature of PINNs are significant advantages. Future works can utilize adaptive collocation point distribution and explore new pre-training and output normalization strategies. The PINN architecture and implementation can be made more computationally efficient to allow for very small time step sizes and training for a higher number of epochs.
@@ Acknowledgments
The authors would like to acknowledge the support from the Office of Naval Research (ONR), USA, to V. Srivastava under grant numbers N00014-21-1-2815 and N00014-21-1-2670.
@@
Asymmetric double-edge notched plate: comparison of PINN (incremental) and PINN (one-time step)
Crack path predictions using PINN (incremental) [1st increment at =0.45 s, time step=0.0025 s, 50000 epochs training] and PINN (one-time step) [=0.65 s, 200000 training epochs] are shown in Figure A1. The collocation grid in Figure 3 was used. Since damage progression is state-dependent, the PINN (incremental) solution exhibits better agreement with the baseline FEM with gradient solution. However, the computationally inexpensive PINN (one-time step) predicts the crack path with reasonable accuracy. Hence, neural network parameter and architecture optimization can be performed in a computationally efficient manner using PINN (one-time step) before final PINN (incremental) applications.
Single-edge notched plate with a hole and plate with four randomly distributed holes: relative error and loss evolution during training
The relative error and loss evolution with the number of epochs during training for both geometries is shown in Figure A2. For the plate with four randomly distributed holes, only is presented due to numerical issues in the interpolation of the relatively complex stress and undamaged free energy fields on the testing grid. The errors decrease monotonically with the number of epochs in this range for both geometries. The losses evolve non-monotonically for both geometries with the number of epochs in this range.
Single-edge notched plate with a hole and plate with four randomly distributed holes: effect of collocation point distribution
Figure A3 shows the different collocation point distributions considered and the corresponding PINN (one-time step) solutions for both geometries. The PINN solutions are approximately insensitive to the point distribution and agree well with the FEM with gradient damage solutions for these two geometries.

### v4

@@ 4.1 Asymmetric double-edge notched plate
A specimen geometry often used to test the predictive capabilities of damage and fracture models in terms of crack path is a plate with asymmetric double-edge notches subjected to tension under plane-strain conditions (Asymmdoublenotch, Asymmdoublenotch1, Asymmdoublenotch2). The loading schematic and specimen geometry considered are shown in Figure 2(A). The crack path predictions obtained using FEM with and without gradient damage for a mesh not aligned with the anticipated crack path are shown in Figure 2(B). As expected, the results from FEM without a gradient damage solution are nonphysical. The significant mesh-dependency of FEM without gradient damage is shown through comparison of the predicted crack paths for two different meshes in Figure 2(C). The nonphysical crack paths are dissimilar for the two meshes. In FEM without gradient damage, upon failure of an element, the subsequent damage growth depends on the local mesh structure around that element, resulting in the solution’s strong mesh dependency.
For the collocation grid in Figure 3, the PINN was applied incrementally [multiple time increments, referred to as PINN (incremental) hereon]. Specifically, the time step size was chosen to be 0.0025 s with the first increment at s. Training for each time increment was performed for 25,000 epochs. The time interval from the initiation of damage to the complete rupture ( = 0.08 s) is small compared to the time to the initiation of damage (FEM with gradient damage: 0.477 s, PINN (incremental): 0.457 s), highlighting the rapid damage growth upon initiation built into the PINN formulation. The crack path predictions from the PINN (incremental) and the baseline FEM with gradient damage solution show good agreement, as shown in Figure 3.
@@ 4.2 Single-edge notched plate with a hole
A single-edge notched plate with a hole under plane-strain conditions is a specimen geometry utilized to validate damage and fracture models (Notchandholemerging, Notchandholemerging1, Notchandholemerging2). The loading schematic and specimen geometry considered are shown in Figure 4(A). The crack path predictions obtained using FEM with and without gradient damage for a mesh not aligned with the anticipated crack path are shown in Figure 4(B). The FEM without gradient damage solution is nonphysical. Figure 4(C) shows the physical solution using FEM without gradient damage for a mesh aligned with the anticipated crack path, highlighting the method’s mesh dependency.
PINN (incremental) is applied to the collocation grid in Figure 5(B) with time step size = 0.01 s, the first time increment at s, and 50,000 epochs training for each time increment. The progression of damage at four different stages and the crack path in the PINN (incremental) solution agree with the FEM with gradient damage solution as seen in Figure 5.
We also consider a variation in this geometry where the hole is moved further away from the notch. Specifically, only the separation of the hole center from the notch centerline in Figure 4(A) is increased by 0.2. The crack path in the FEM with gradient damage solution changes significantly as seen in Figure 6. The PINN [ s, time step size = 0.55 s, 200,000 epochs training for the time increment] captures the crack path variation with the defect positioning accurately, as shown in Figure 6.
@@ 4.3 Plate with four randomly distributed holes
We now consider a relatively complex geometry consisting of four holes with different diameters randomly distributed on the 1x1 plate domain. The loading schematic, specimen geometry, and finite element mesh considered are shown in Figure 7(A). The crack path prediction obtained using FEM with gradient damage is shown in Figure 7(B).
The PINN (incremental) is applied to the collocation grid in Figure 8(B) with time step size = 0.0025 s, the first time increment at s, and 25,000 epochs training for each time increment. Good agreement of the damage progression at four different stages and the crack path from the PINN (incremental) solution with the baseline FEM with gradient damage solution can be seen in Figure 8.
@@ 5 Assessment of the PINN performance for fracture
Previously, we demonstrated the success of PINN without gradient damage formulation in predicting the fracture path for a variety of problems. The performance of any new and emerging computational method has to be evaluated to guide the choice of underlying parameters for future applications. Towards this, we investigate the effect of a few key neural network parameters on the PINN’s prediction performance. The PINN solutions in this section for the asymmetric double-edge notched plate geometry are obtained using a relatively computationally inexpensive single time increment approach [referred to as PINN (one-time step) hereon]. In this work, we have considered (i) hyperelastic (rate-independent) material response, (ii) instantaneous damage growth, which is approximately observed in experiments for elastomers and several soft polymers, where damage, when it initiates, grows very rapidly, and (iii) monotonic loading. Under these conditions, the PINN (one-time step) solutions provide crack paths with reasonable accuracy, as shown in Appendix Figure A3. For the general class of fracture problems, PINN (incremental) [multiple time increments] solutions will have better accuracy. However, PINN (one-time step) can enable computationally cheaper neural network parameter and architecture optimization before PINN (incremental) is applied for accurate final results.
The crack paths are obtained through PINN (one-time step) solutions at s with 100,000 epochs of training. relative errors and losses are evaluated through PINN (one-time step) application at s (before damage initiation) with 100,000 epochs training. First, variations in the collocation point distribution (approximately constant grid density) are considered. Further, the number of layers, the number of neurons in each layer, and the grid density (in a fixed grid structure) are systematically varied. The base number of layers, number of neurons in each layer, and grid density are the values used in Section 4, i.e., , , and , respectively. We also compare the computation time for the PINN using the base network and base grid density with those for FEM without and with gradient damage. The effects of simple pre-training and output normalization strategies on the solutions are also explored.
@@
relative error ( - displacement, - stress, UFE - undamaged free energy density) and loss (total, - balance of linear momentum, - boundary conditions, - constitutive equations) evolution with number of epochs during training [PINN (one-time step), =0.45 s, 100,000 epochs, grid from Figure 3] for the asymmetric double-edge notched plate geometry is shown in Figure 9. The errors and losses decrease monotonically with the number of epochs in this range. Corresponding results for the single-edge notched plate with a hole and the plate with four randomly distributed holes geometries are presented in the Appendix. The low errors with respect to the baseline FEM solutions highlight the validity of the simplification discussed in Section 3.1.1.
@@ 5.3 Effect of the number of layers, number of neurons in each layer, and grid density
The three different numbers of layers, number of neurons in each layer, and grid density considered were 3, 5, 8; 10, 30, 50; and , , , respectively. A numerical Factorial Design of Experiments (DOE) was performed to understand the main, simple, and interacting effects of the three hyperparameters. The complete results for the relative errors and the crack paths for the DOE study are presented in the Appendix. Figures 11(A), (B), (C) and (D) show the % errors for a few cases, highlighting the effect of (i) number of layers ( grid density), (ii) number of neurons in each layer ( grid density),(iii) grid density (5 layers), and (iv) grid density (30 neurons in each layer), respectively. All three factors have main effects. Significant interaction effects are observed between the number of layers and the number of neurons, as well as between the grid density and the number of layers. Underfitting or overfitting occurs when the neural network parameters are not optimal (see Figures A4 and A5).
5.4 Comparison of computation time: PINN, FEM without and with gradient damage
The computation time using the base network (5 layers each with 30 neurons) and the base grid density () for 1 epoch of PINN training remained approximately constant for all numerical problems considered and was equal to 0.06 s (GPU accelerated, QuadroRTX, CCV@Brown), 0.18 s (GPU accelerated, NVIDIA T600 & Intel Core i7-10700 Processor, Dell Precision 3650 Tower), and 0.24 s (not GPU accelerated, CCV@Brown). FEM without and with gradient damage (both implemented in ABAQUS/Explicit (2018)) were parallelized using 32 processors on a system with an Intel Xeon Gold 6326 CPU. For the finite element meshes considered in this work (10,000 elements), the approximate computation time for an increment with a time step size of approximately s was 0.003 s for both FEM approaches.
PINN (incremental) in general requires more computation time than FEM. The complexity of computationally implementing gradient damage models in FEM and the associated implementation time costs are important considerations. Some reduction in computation time for PINN (incremental) can be achieved when the material response is rate-independent. As done in this work for an elastomer with hyperelastic response, the solution evaluation in PINN (incremental) for a rate-independent material can be started at any time instant before damage initiation. The computation time for PINN (one-time step), however, can be comparable to FEM.
@@ 5.5 Effect of pre-training
Pre-training has been shown in the literature to improve the accuracy and convergence of PINNs (Pretrainingone, Pretrainingtwo). We investigate the effect of one pre-training strategy. Specifically, the domain seeded with a uniform 80 x 80 collocation point grid is assigned the same material properties and boundary conditions discussed in Section 4. The trained network parameters from stretching this unit square ( = 0.45 s in one-time step, 100,000 epochs) are used to initialize the PINN for application to the asymmetric double-edge notch plate geometry. Comparison between the PINN (one-time step) results (at = 0.45 s [before damage initiation] and 0.65 s [after damage initiation], 100,000 epochs) with and without pre-training in Figure 12 shows the simple pre-training associated convergence and accuracy enhancement.
@@ 5.6 Effect of PINN output normalization
Normalization of PINN outputs is another technique that has been shown in the literature to enhance accuracy and convergence (Outputnormalization1, Outputnormalization2). We consider a simple normalization scheme in the form of fixed normalization numbers that ensures physical consistency. Specifically, are normalized by , respectively before evaluating the loss functions in equations (3.11), (3.13), (3.14). Figure 13 shows the comparison between the PINN (one-time step) results (at = 0.45 s [before damage initiation] - , , , 100,000 epochs and 0.65 s [after damage initiation] - , , , 100,000 epochs) for the asymmetric double-edge notched plate geometry with and without output normalization. Faster convergence and increased accuracy are achieved with this simple output normalization scheme.
@@ 6 Conclusions
The important and evolving field of fracture modeling commonly uses the finite element method (FEM). FEM without gradient damage is relatively easy to implement. However, the corresponding results are highly mesh-dependent and often inaccurate. FEM with gradient damage enables mesh-independent predictions at the cost of increased mathematical and numerical implementation complexities. Specifically, additional spatial gradient calculations and complex user-defined element and user subroutine development are necessary. Physics-informed neural networks (PINNs) have recently emerged as a computational method that can be applied to solve solid mechanics problems. The limited studies on PINNs for damage and fracture modeling in the literature are confined to the small-strain regime and use the gradient damage formulation by default. Due to the meshless nature of PINNs, we posed the question of whether we can correctly solve fracture problems without needing the gradient damage formulation. We propose a PINN without gradient damage to model large deformation fracture in elastomers. The PINN does not require training data and utilizes the collocation method. The capability of the PINN to accurately predict the crack path has been validated by studying a variety of benchmark defect configuration problems, where we used accurate solutions from FEM with gradient damage and a highly refined mesh. We also conducted an assessment of the PINN as a fracture modeling method by systematically varying a few key parameters to guide future design of PINNs for fracture problems. The PINN crack path predictions are largely insensitive to the collocation point distribution. The proposed modeling framework and neural network architecture could motivate the application of PINNs without gradient damage for modeling fracture in a broad class of materials.
The relatively new nature of PINNs as a numerical method implies that the solution quality depends on a range of factors, and problem-dependent insights are necessary for the successful application of PINNs. However, the relative ease of mathematical formulation, numerical implementation, and mesh-free nature of PINNs are significant advantages. Future work can utilize adaptive collocation point distribution and explore new pre-training and output normalization strategies. The PINN architecture and implementation can be made more computationally efficient to allow for very small time step sizes and training for a higher number of epochs.
@@ Acknowledgments
The authors thank Dr. Khemraj Shukla at Brown University for discussions on the topic. The authors would like to acknowledge and thank the support from the Office of Naval Research (ONR), USA, to V. Srivastava under grants N00014-21-1-2815 and N00014-23-1-2688.
@@
Asymmetric double-edge notched plate
Effect of loss weights
The individual loss terms (using all weights = 1) for the asymmetric double-edge notched plate ( s, 100,000 training epochs) are not in the same order of magnitude as seen in Figure 9. Particularly, the constitutive loss is significantly higher than the balance of linear momentum (BLM) and the boundary condition (BC) losses. The loss weights can be adjusted to further minimize the total loss. First, only was increased (= 10, 100) to reduce . Figure A1(A) shows the corresponding loss and error evolution with training epochs. decreased with increasing as anticipated. However, both and increased, with a significant increase in . The total loss also increased along with the errors. Since was the most sensitive to an increase in , we also simultaneously increased ( = = 10 and 100). A similar observation was made with increasing with the weights, and hence the total loss increasing as seen in Figure A1(B). The errors changed insignificantly compared to the baseline (all weights = 1) in this case. Finally, the three weights were simultaneously increased ( = = = 10 and 100). The corresponding losses and error evolution were similar to the baseline, as shown in A1(C). Hence, it was concluded that the baseline weight values provide a good balance between loss and error minimization. It can be noted that the optimal loss weight values can vary for different sets of material parameters (e.g., elastic modulus orders of magnitude different from that used in this study). Suitable weights need to be obtained for specific boundary value and material type problems.
PINN solution convergence
The convergence of the PINN solutions for the asymmetric double-edge notched plate geometry is verified with respect to the number of training epochs. The error, loss ( s), and final crack path ( s) evolution with training epochs (1,000,000 epochs - 10 times the number of epochs used) is shown in Figure A2. The errors and losses decrease monotonically and approximately converge. The overall width of the crack path decreases with epochs, and the path converges. The PINN formulation does not have an inherent length scale parameter to control the damage zone width. Consequently, the predictions are a function of the neural network parameters (see Figure A5). There is no physical correlation. Even in the baseline FEM with gradient damage approach, where a diffusive crack is considered, the “length scale” parameter is present purely for numerical reasons.
Comparison of PINN (incremental) and PINN (one-time step)
The crack path predictions using PINN (incremental) [1st increment at =0.45 s, time step=0.0025 s, 50,000 epochs training] and PINN (one-time step) [=0.65 s, 100,000 training epochs] are shown in Figure A3. The PINN (incremental) crack path solution exhibits better agreement with the baseline FEM with gradient solution. However, the computationally inexpensive PINN (one-time step) predicts the crack path with reasonable accuracy. Hence, neural network parameter and architecture optimization can be performed in a computationally efficient manner using PINN (one-time step) before PINN (incremental) applications for accurate final results.
@@
Single-edge notched plate with a hole and plate with four randomly distributed holes
relative error and loss evolution during training
The relative error and loss evolution with the number of epochs during training for both geometries is shown in Figure A6. For the plate with four randomly distributed holes, only is presented due to numerical issues in the interpolation of the relatively complex stress and undamaged free energy density fields on the testing grid. The errors and losses decrease monotonically with the number of epochs in this range for both geometries.
Effect of collocation point distribution
Figure A7 shows the different collocation point distributions considered and the corresponding PINN (one-time step) crack path solutions for both geometries. The PINN solutions are approximately insensitive to the collocation point distribution and agree well with the baseline FEM with gradient damage solutions for both geometries.